# Memento-II: Learning by Stateful Reflective Memory

### v1

@@ 6 Conclusions
Looking ahead, there are several promising directions for future work. These include extending the analysis to continuous state spaces, developing more advanced strategies for memory curation, and conducting large-scale empirical validation on complex domains. Together, these efforts could further strengthen the role of reflection and memory in advancing memory-augmented LLM agents.
@@ 7 Acknowledgements
We are grateful to Xue Yuan for valuable discussions on reinforcement learning and kernel methods, Siyuan for insights into bandit algorithms and the discovery–exploration–exploitation trade-off, and Huichi and Yihang for assistance with implementation. We also thank collaborators who reviewed earlier versions of this manuscript and provided helpful feedback.
@@ Appendix B Proof of Theorem 7
Unlike standard soft reinforcement learning [Haarnoja2018], our setting involves state-dependent base measures that vary with memory content, requiring a careful analysis of the induced operators. In particular, extending KL-regularised evaluation requires establishing that the associated log-sum-exp operations preserve contraction for both the value function and the prior distributed resulted from memory updates. Lemma 16 extends the log-sum-exp Lipschitz bound [BoydVandenberghe2004] to this setting with Parzen base measures and memory updates.
@@ B.1 Bi-Lipschitz Properties
$f(q)\;\triangleq\;f(b,q)=\alpha\log\sum_{i=1}^{n}b_{i}\exp\!\left(\frac{q_{i}}{\alpha}\right).$
@@ B.2 Contraction of the KL-Regularised Evaluation Operator
$(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)\;\triangleq\;r(x,c)+\gamma\,\alpha\log\!\Big(\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})\exp(Q(x^{\prime},c^{\prime})/\alpha)\Big),$ (23)
@@
$f(b,q)\triangleq\alpha\log\sum_{i}b_{i}e^{q_{i}/\alpha},\qquad q\triangleq Q(x^{\prime},\cdot),\quad q^{\prime}\triangleq Q^{\prime}(x^{\prime},\cdot).$
@@ B.4 Convergence of Read-Write Iterations
Let $(Q_{k},\mu_{k})$ be the sequence of iterates. By Theorem 19, the soft values satisfy $V^{\mu_{k+1}}(x)\geq V^{\mu_{k}}(x)$ for all $x$.
@@
The last term uses $\mathrm{KL}(\nu\|\mu_{0})\geq-\log|M|$ for any $\nu\in\Delta(M)$. Since $\{V^{\mu_{k}}\}$ is bounded and monotone increasing, it converges pointwise to some limit $\bar{V}$. By Theorem 17, each evaluation stage converges to the unique $Q^{\mu_{k}}$. The sequence $\{Q^{\mu_{k}}\}$ lies in a compact subset of bounded functions. Any limit point $(\bar{Q},\bar{\mu})$ must satisfy:
@@ Appendix C Lipschitz Properties of Memory Updates
$\|M^{\prime}-M\|\;\triangleq\;\sup_{x}\mathrm{TV}\!\left(\mu_{0}^{M^{\prime}}(\cdot\mid x),\mu_{0}^{M}(\cdot\mid x)\right)\;=\;\frac{1}{2}\sup_{x}\big\|\mu_{0}^{M^{\prime}}(\cdot\mid x)-\mu_{0}^{M}(\cdot\mid x)\big\|_{1}.$ (33)
@@
For all $x\in\mathcal{X}$ and $c\in\mathcal{C}$, $Q(x,c)\in[Q_{\min},Q_{\max}]$ and $\Delta_{Q}\triangleq Q_{\max}-Q_{\min}$.
@@
$G_{M}(x;Q)\;\triangleq\;\alpha\log\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp\!\big(Q(x,c^{\prime})/\alpha\big),$
@@
$\Phi_{M}(c\mid x;Q)\;\triangleq\;\frac{\mu_{0}^{M}(c\mid x)\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp(Q(x,c^{\prime})/\alpha)}.$
@@ D.3 Slow Timescale: Memory Dynamics
with $\varsigma$ denoting slow time. By Assumption 8(iv), this ODE admits a compact attractor set $\mathcal{M}_{\infty}$.
@@ D.6 Convergence to the Limit Set
By Lemma 26, $\epsilon_{t}\to 0$ almost surely. Since $M_{t}\to\mathcal{M}_{\infty}$, all limit points of $(Z_{t},M_{t})$ lie in
$\mathcal{L}=\{(Z_{M}^{\star},M):M\in\mathcal{M}_{\infty}\}.$
@@
Appendix E Comparison between SRDP and Reflected MDP
The Stateful Reflection Decision Process (SRDP; Definition 2) specifies the generative mechanism by which an LLM-based agent interacts with its environment through episodic memory, reflection, and action execution. It explicitly models the internal read–write structure of reflection and the stochastic action selection mediated by the LLM.
The Reflected MDP (Definition 3) is an induced control process obtained by collapsing the internal reflection–action pipeline of SRDP into a single Markovian decision layer over augmented states. It captures the effective dynamics perceived by the agent after the reflection mechanism has been fixed, and serves as the appropriate abstraction for theoretical analysis and reinforcement learning.
Table 1 summarises the key distinctions between the two formulations.
Table 1: Comparison between SRDP and Reflected MDP.
Aspect SRDP Reflected MDP
Role in the framework Generative decision process describing how reflection, memory, and actions are produced. Induced Markov decision process capturing the effective control dynamics.
Definition level Explicitly models internal reflection (read/write) and LLM-mediated action sampling. Abstracts away internal mechanisms into a single-stage decision process.
State space $\mathcal{S}$: environment states, with memory handled as an external component. $\mathcal{X}=\mathcal{S}\times\mathcal{M}$: augmented states including memory.
Action space $\mathcal{A}$: primitive environment actions produced by the LLM. $\mathcal{C}$: reflection (retrieval) actions that index memory cases.
Decision mechanism Two-stage: retrieval policy $\mu(s,M)$ followed by LLM kernel $p_{\mathrm{LLM}}(a\mid s,c)$. Single-stage: choosing $c$ directly determines expected transitions and rewards.
Transition kernel Environment dynamics $\mathcal{P}(s^{\prime}\mid s,a)$ applied after LLM action generation. $\widetilde{\mathcal{P}}(x^{\prime}|x,c)=\mathbb{E}_{a\sim p_{\mathrm{LLM}}(\cdot|s,c)}\big[\mathcal{P}(s^{\prime}\mid s,a)\,\delta_{M^{\prime}}(\mathsf{Write}())\big].$
Reward function $\mathcal{R}(s,a)$. $\widetilde{\mathcal{R}}(x,c)=\mathbb{E}_{a\sim p_{\mathrm{LLM}}(\cdot\mid s,c)}[\mathcal{R}(s,a)]$.
Memory dynamics Memory is explicitly read and written as part of the interaction loop. Memory evolution is embedded into state transitions over $\mathcal{X}$.
Markov property Not Markovian in $\mathcal{S}$ due to latent reflection and memory updates. Markovian in the augmented state space $\mathcal{X}$.
Interpretation Reflective process: how the agent reasons and acts. Reflected process: how the environment appears after reflection is folded in.
Primary use Conceptual and mechanistic description of LLM-based reflection. Theoretical analysis, planning, and reinforcement learning.

### v2

@@ 6 Conclusions and Future Work
Looking ahead, several promising directions for future work emerge. In particular, we plan to internalise episodic memory within the LLM architecture itself, yielding a unified stateful neural machine, which we outline next.
6.1 Internalising Episodic Memory into the LLM Architecture
In modern contexts, LLM agents increasingly employ explicit episodic memory systems to extend context and improve performance across diverse tasks [Zhong2024, Chhikara2025, fountas2024human]. Yet, the integration of memory retrieval with decision-making in LLMs still lacks a firm theoretical foundation, particularly regarding the conditions under which such systems converge to optimal behaviour. Our work addresses this gap directly.
A natural next step is to integrate the episodic memory and read–write reflection mechanism of SRDP directly into the LLM architecture, yielding a single neural system whose internal state corresponds to the augmented state $(s,M)$. Recent “test-time memorisation” lines of work provide useful architectural blueprints. In particular, Titans introduces a neural long-term memory module that updates while the model is running, thereby supporting continual memorisation beyond the context window without offline retraining [behrouz2024titans]. MIRAS further reframes modern sequence models as instances of associative memory with explicit choices of memory structure, retention, and online optimisation [behrouz2025miras]. Our EM-LLM work in [fountas2024human] introduces a human-inspired episodic memory architecture that enables large language models to operate with effectively infinite context by selectively storing, retrieving, and reusing past experiences to support long-horizon reasoning and decision-making. These developments suggest a promising direction: replace external retrieval of SRDP over a growing case base by an internal memory state that is updated online, while retaining the SRDP control view in which “read” implements policy improvement and “write” implements policy evaluation.
Our Read-Write loop differs from Titans-style test-time updates in a key way: SRDP performs stateful adaptation through experience-grounded reflection and memory evolution, rather than gradient-based parameter updates of the base model. A concrete research direction is to design a hybrid architecture in which the LLM remains fixed or slow-updated, but an internal memory module (or a small retrieval policy head) is updated online using environment feedback, with a provable two-time-scale separation between fast policy iteration and slow memory evolution. This would preserve the stability benefits of our reflected-MDP analysis while inheriting the efficiency and scalability motivations emphasised in Titans/MIRAS [behrouz2024titans, behrouz2025miras] and EM-LLM [fountas2024human].
More broadly, recent work on test-time training layers and nested optimisation provides a conceptual bridge between architecture and learning dynamics. Test-time training (TTT) instantiates sequence layers whose hidden state is itself a learnable model updated during inference [sun2024ttt]. Nested Learning argues that modern deep models can be viewed as systems of nested optimisation problems with their own internal workflows, offering a principled lens on continual learning and catastrophic forgetting [behrouz2025nested]. An important future direction is to recast SRDP within this nested-optimisation view: the outer loop corresponds to fixed LLM representations and long-term objectives, while the inner loop implements read–write reflective adaptation as an online optimisation over an internal memory state. Such a unification could clarify when reflection-based memory updates are sufficient for continual learning, when parameter adaptation is necessary, and how to combine both without destabilising long-horizon control.
Another dimension worth exploring is principled memory consolidation. Recent work on KV-state consolidation in Transformer architectures (Bottlenecked Transformers) demonstrates that selectively rewriting and compressing internal memory can improve predictive efficiency and long-horizon reasoning without expanding context length [oomerjee2025bottlenecked]. From the perspective of SRDP, such mechanisms provide an architectural realisation of the write operation, transforming episodic experience into a compact latent state that supports future reflection. This view is further supported by our work on Agent K, which shows that structured cycles of interaction, reflection, abstraction, and consolidation enable effective experiential learning without parameter fine-tuning [Grosnit2024KolbBasedEL]. Together, these results point toward an architecture in which episodic memory is consolidated on a slow time scale and queried on a fast time scale, aligning architectural dynamics with the SRDP formalism.
Finally, these connections open new theoretical questions. MIRAS links sequence modelling to online optimisation and retention mechanisms [behrouz2025miras]; integrating SRDP into this framework may yield sharper characterisations of memory coverage, approximation error, and the conditions under which reflective memory updates guarantee monotone improvement and asymptotic optimality. Establishing such results for neural internal memories (rather than external case stores) would move SRDP toward a a unified neural state machine architecture in which episodic experience, reflection, and control are implemented end-to-end together.
@@ 7 Acknowledgements
We are grateful to Xue Yuan for valuable discussions on reinforcement learning and kernel methods, Siyuan Gu for insights into bandit algorithms and the discovery–exploration–exploitation trade-off, and Huichi Zhou and Yihang Chen and Yuxiang Chen for assistance with implementation and proof read. We also thank collaborators who reviewed earlier versions of this manuscript and provided helpful feedback.
@@ Appendix B Proof of Theorem 8
Unlike standard soft reinforcement learning [haarnoja2018soft], our setting involves state-dependent base measures that vary with memory content, requiring a careful analysis of the induced operators. In particular, extending KL-regularised evaluation requires establishing that the associated log-sum-exp operations preserve contraction for both the value function and the prior distributed resulted from memory updates. Lemma 16 extends the log-sum-exp Lipschitz bound [BoydVandenberghe2004] to this setting with Parzen base measures and memory updates.
@@ B.1 Bi-Lipschitz Properties
$f(q)\;:=\;f(b,q)=\alpha\log\sum_{i=1}^{n}b_{i}\exp\!\left(\frac{q_{i}}{\alpha}\right).$
@@ B.2 Contraction of the KL-Regularised Evaluation Operator
$(\mathcal{T}^{\mu,M}_{\mathrm{KL}}Q)(x,c)\;:=\;r(x,c)+\gamma\,\alpha\log\!\Big(\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x^{\prime})\exp(Q(x^{\prime},c^{\prime})/\alpha)\Big),$ (21)
@@
$f(b,q):=\alpha\log\sum_{i}b_{i}e^{q_{i}/\alpha},\qquad q:=Q(x^{\prime},\cdot),\quad q^{\prime}:=Q^{\prime}(x^{\prime},\cdot).$
@@ B.4 Convergence of Read-Write Iterations
Let $(Q_{t},\mu_{t})$ be the sequence of iterates. By Theorem 19, the soft values satisfy $V^{\mu_{t+1}}(x)\geq V^{\mu_{t}}(x)$ for all $x$.
@@
The last term uses $\mathrm{KL}(\nu\|\mu_{0})\geq-\log|M|$ for any $\nu\in\Delta(M)$. Since $\{V^{\mu_{t}}\}$ is bounded and monotone increasing, it converges pointwise to some limit $\bar{V}$. By Theorem 17, each evaluation stage converges to the unique $Q^{\mu_{t}}$. The sequence $\{Q^{\mu_{t}}\}$ lies in a compact subset of bounded functions. Any limit point $(\bar{Q},\bar{\mu})$ must satisfy:
@@ Appendix C Lipschitz Properties of Memory Updates
$\|M^{\prime}-M\|\;:=\;\sup_{x}\mathrm{TV}\!\left(\mu_{0}^{M^{\prime}}(\cdot\mid x),\mu_{0}^{M}(\cdot\mid x)\right)\;=\;\frac{1}{2}\sup_{x}\big\|\mu_{0}^{M^{\prime}}(\cdot\mid x)-\mu_{0}^{M}(\cdot\mid x)\big\|_{1}.$ (31)
@@
For all $x\in\mathcal{X}$ and $c\in\mathcal{C}$, $Q(x,c)\in[Q_{\min},Q_{\max}]$ and $\Delta_{Q}:=Q_{\max}-Q_{\min}$.
@@
$G_{M}(x;Q)\;:=\;\alpha\log\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp\!\big(Q(x,c^{\prime})/\alpha\big),$
@@
$\Phi_{M}(c\mid x;Q)\;:=\;\frac{\mu_{0}^{M}(c\mid x)\exp(Q(x,c)/\alpha)}{\sum_{c^{\prime}\in\mathcal{C}}\mu_{0}^{M}(c^{\prime}\mid x)\exp(Q(x,c^{\prime})/\alpha)}.$
@@ D.3 Slow Timescale: Memory Dynamics
with $\varsigma$ denoting slow time. By Assumption 9(iv), this ODE admits a compact attractor set $M_{\infty}$.
@@ D.6 Convergence to the Limit Set
By Lemma 26, $\epsilon_{t}\to 0$ almost surely. Since $M_{t}\to M_{\infty}$, all limit points of $(Z_{t},M_{t})$ lie in
$\mathcal{L}=\{(Z_{M}^{\star},M):M\in M_{\infty}\}.$
@@
Appendix E Notations of MDP, SRDP, and Reflected MDP
Table 1 summarises and contrasts the notation used for the underlying Markov Decision Process (MDP), the Stateful Reflective Decision Process (SRDP) (Definition 3), and its Markovian reformulation as the Reflected MDP (Definition 4). The underlying MDP models only the environment dynamics, abstracting away any internal structure of the agent. The SRDP extends this framework by explicitly modelling episodic memory and a two-stage decision mechanism, in which a retrieval action selects a memory case that conditions the LLM’s action generation. Since the evolving memory breaks the Markov property in the original state space, the Reflected MDP restores Markovianity by augmenting the state with memory and absorbing the fixed LLM behaviour into effective transition and reward kernels. This reformulation allows classical reinforcement learning analysis to be applied while preserving the full agent–memory–LLM interaction dynamics.
Concept Underlying MDP SRDP Reflected MDP
Decision process $\mathcal{D}=\langle\mathcal{S},\mathcal{A},P,R,\gamma\rangle$ $\mathcal{D}_{\mathrm{SRDP}}=\langle\mathcal{S},\mathcal{A},P,R,\gamma,\mathcal{M},p_{\mathrm{LLM}}\rangle$ $\mathcal{D}_{\mathrm{ReMDP}}=\langle\mathcal{X},\mathcal{C},P_{\mathrm{LLM}},R_{\mathrm{LLM}},\gamma\rangle$
State space $s\in\mathcal{S}$ $s\in\mathcal{S}$ (environment state) $x=(s,M)\in\mathcal{X}:=\mathcal{S}\times\mathcal{M}$
Action space $a\in\mathcal{A}$ Two-stage: retrieval $c\in\mathcal{C}(M)$, then $a\in\mathcal{A}$ Retrieval only: $c\in\mathcal{C}(M)$
Memory None Episodic memory $M\in\mathcal{M}$, evolving via $\mathrm{Write}$ Included in state: $M$ is part of $x=(s,M)$
Retrieval action
set
N/A $\mathcal{C}(M):=M$ (memory indexers) $\mathcal{C}(M):=M$
Policy $\pi(a\mid s)$ Composite policy: $\pi^{\mu}(a\mid s,M)=\sum_{c\in M}\mu(c\mid s,M)\,p_{\mathrm{LLM}}(a\mid s,c)$ Retrieval policy $\mu(c\mid x)$
LLM kernel N/A $p_{\mathrm{LLM}}(a\mid s,c)$ Absorbed into $P_{\mathrm{LLM}},R_{\mathrm{LLM}}$
Transition kernel $P(s^{\prime}\mid s,a)$ $s^{\prime}\sim P(\cdot\mid s,a)$,
$M^{\prime}=\mathrm{Write}(M,s,a,r,s^{\prime})$ $P_{\mathrm{LLM}}(x^{\prime}\mid x,c)=\sum_{a}p_{\mathrm{LLM}}(a\mid s,c)\,P(s^{\prime}\mid s,a)$
Reward $R(s,a)$ $r=R(s,a)$ $R_{\mathrm{LLM}}(x,c)=\sum_{a}p_{\mathrm{LLM}}(a\mid s,c)R(s,a)$
Markov property Markov in $s$ Non-Markov in $s$ due to evolving $M$ Markov in augmented state $x=(s,M)$
Learning focus Action selection Retrieval + memory growth Retrieval policy optimisation
Role Models environment
dynamics
Models agent–memory–LLM interaction Markovian control abstraction of SRDP
Table 1: Comparison of notations for the underlying MDP, Stateful Reflective Decision Process (SRDP), and the Reflected MDP.